# Surface-mediated reduction of radiation damage in tungsten revealed by advanced ion channeling analysis

### v1

@@
This work was partially carried out within the framework of the EUROfusion Consortium, funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No 101052200 — EUROfusion). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them. The authors wish to acknowledge CSC – IT Center for Science, Finland, for computational resources, and the Finnish Computing Competence Infrastructure (FCCI) for supporting this project with computational and data storage resources. Xin Jin thanks to Shenzhen Science and Technology Program (ZDSYS20230626091501002).

### v2

@@ 4 Conclusions
Acknowledgements
The authors wish to acknowledge CSC – IT Center for Science, Finland, for computational resources, and the Finnish Computing Competence Infrastructure (FCCI) for supporting this project with computational and data storage resources. Xin Jin thanks to Shenzhen Science and Technology Program (ZDSYS20230626091501002).
Funding
This work was partially carried out within the framework of the EUROfusion Consortium, funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No 101052200 — EUROfusion). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them. Xin Jin thanks to Shenzhen Science and Technology Program (ZDSYS20230626091501002) for financial support.